OpenStudy (anonymous):

find \[\frac{ \delta z }{ \delta x }\]and\[\frac{ \delta z }{ \delta y }\]where\[\frac{ yz }{ x }-\sin x \cos^{2} y =ye^{yz}-z(xy+1)^{3} \]\[z = f(x,y)\]

OpenStudy (unklerhaukus):

Can you take the partial derivative of the first term with respect to x?

OpenStudy (anonymous):

erm, do u mean del/del x (yz/x)?

OpenStudy (unklerhaukus):

yesh

OpenStudy (anonymous):

ya, no prob wait, i write the eqn

OpenStudy (anonymous):

del/del x (yzx^(-1))=(del z/del x )yx^(-1) + (-1)yzx^(-2)

OpenStudy (anonymous):

= (del z / del x)(y/x) -yz/(x^(2)) right?

OpenStudy (unklerhaukus):

yeah that is correct, good work , now , the other terms

OpenStudy (anonymous):

(del / del x )(sin (x) cos ^2(y)) = cos (x) cos^2 (y)

OpenStudy (perl):

what does Del mean here

OpenStudy (perl):

, delta ?

OpenStudy (unklerhaukus):

∂, partial derivative \(\partial \) \partial

OpenStudy (perl):

ok thanks

OpenStudy (unklerhaukus):

once you have take the implicit partial derivative of the both sides of the equation, you just have to rearrange to solve for \(\dfrac{\partial z}{\partial x }\)

OpenStudy (anonymous):

(del / del x)ye^(yz)= (ye^(yz))((del z/ del x)(y))=(del z/ del x)y^2e^(yz)

OpenStudy (anonymous):

i still hv some problem, that is my ans and ans in wolframalpha is different, is that my ans wrong?? i'll show all steps... but i cant find any exponent in wolframalpha

OpenStudy (unklerhaukus):

maybe wolfram doesn't know that z(x,y)

OpenStudy (anonymous):

i dont think my solution can be simplify some more... but wolfram seem like have some way to simplify

OpenStudy (perl):

the original equation is y*x^(-1)*z - sin x * cos^2(y) = y*e^(y*z)- z(xy+1)^3

OpenStudy (anonymous):

ya

OpenStudy (perl):

partial z with respect to x (implicitly is) I will use z'_x for partial z with respect to x y(-1)x^(-2)z + y(x^(-1)z'_x = y*e^(y*z)*y*z'_x - z'_x(xy + 1)^3 -z(3)(xy+1)^2*y

OpenStudy (unklerhaukus):

@neoh147 , i got the same as you on that .png \[\frac{\partial z}{\partial x}=z_x=\dfrac{\frac{yz}{x^2}+\cos x\cos^2y-3zy(xy+1)^2}{\frac yx-y^2e^{yz}+(xy+1)^3}\]

OpenStudy (perl):

I can do it in maple, one moment

OpenStudy (anonymous):

ok, that means probably my calculation is nothing wrong... can u help me to check the partial diff respect to y??

OpenStudy (perl):

I can give you maple solution, how do upload a .png?

OpenStudy (anonymous):

use print screen, paste it on paint and attach file

OpenStudy (perl):

one sec

OpenStudy (anonymous):

ok ^_^

OpenStudy (perl):

its ugly though ;)

OpenStudy (anonymous):

ok, but my answer is require in the simplest form, not expanded form, anyway thx for sharing the eqn ^_^

OpenStudy (unklerhaukus):

@neoh147 the third line on your ∂/∂y have yo got the right sign on your derivative of that trig function?

OpenStudy (anonymous):

ya, suppose differentiate cos y should be - sin y, I make a careless mistake ><

OpenStudy (unklerhaukus):

and i think you missed a \(y\) also \[\frac{\partial z}{\partial y}(yx^{-1}-y^{\color{red}2}e^{yz}+...\]

OpenStudy (anonymous):

ok >w<

OpenStudy (unklerhaukus):

going back to the first one , if you want to simplify\[\frac{\partial z}{\partial x}=\dfrac{\frac{yz}{x^2}+\cos x\cos^2y-3zy(xy+1)^2}{\frac yx-y^2e^{yz}+(xy+1)^3}\] multiply numerator and denominator by x^2 \[=\dfrac{yz+x^2\cos x\cos^2y-3x^2yz(xy+1)^2}{xy-x^2y^2e^{yz}+x^2 (xy+1)^3}\]

OpenStudy (unklerhaukus):

@perl can you ask maple to factorise your results,?

OpenStudy (perl):

let me try just type factor?

OpenStudy (perl):

wont let me

OpenStudy (anonymous):

ok, i get it d, the complex fractions must be reduce to a single fraction right?

OpenStudy (unklerhaukus):

i don't know maple @perl , but i think that you can factor you results they will look much more similar to the results that have been derived manually

OpenStudy (unklerhaukus):

i dont know how to simplify any further

OpenStudy (anonymous):

ok thx a lot @UnkleRhaukus and @perl ^_^

OpenStudy (perl):

partial z / partial x looks right

OpenStudy (perl):

maple just factored out one x from the denominator, i guess the only common term

OpenStudy (perl):